a rectangle has the perimeter of 58in. the length is 5in longer than the width. Find the Dimensions of the rectangle.
Answer by ewatrrr(24785)

 
Hi,
.
L + w = 29
(w+5) + w = 29
2w = 24
w = 12in and L: = 17in
And...checking
L + w = 29